OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2)

OpenStudy (anonymous):

slope is: \[\frac{ y_2 - y_1 }{ x_2 - x_1 }\]

OpenStudy (anonymous):

standard form is: \[y - y_1 = m * (x - x_1)\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

lol medal me up

OpenStudy (anonymous):

i think i did lol i have never used this before